A new system of unit is choosen in which the unit of mass is taken 10kg, unit of length is taken 100m and the unit of time is taken 100s. What will be the unit of speed of light in this new system of units?

Grade:11

1 Answers

Arun
6 years ago
Speed of light in original system = 3 *108 m/sec
Now 
According to new system
Unit of length = 100 m
Unit of time = 100 sec
Hence in new system
Speed of light = 3*108 *100/100 2 = 
= 3* 106